### v1

# Machine Learning and First-Principles Predictions of Materials with Low Lattice Thermal Conductivity

Abstract
We perform machine learning (ML) simulations and density functional theory (DFT) calculations to search for materials with low lattice thermal conductivity, $\kappa_{L}$. Several cadmium (Cd) compounds containing elements from the alkali-metal and carbon groups including A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) are predicted by our ML models to exhibit very low $\kappa_{L}$ values ($<1.0$ W/mK), rendering these materials suitable for potential thermal management and insulation applications. Further DFT calculations of electronic and transport properties indicate that the figure of merit, $ZT$, for thermoelectric performance can exceed 1.0 in compounds such as K2CdPb, K2CdSn, and K2CdGe, which are thereby also promising thermoelectric materials.
lattice thermal conductivity, thermoelectric material, machine learning, density functional theory
I Introduction
Materials with low lattice thermal conductivity ($\kappa_{L}$) have important applications in thermal management and energy conversion, by serving as thermal insulation and barrier coatings, or as thermoelectrics. In particular, thermoelectric (TE) materials can directly convert between thermal and electrical energy [1, 2, 3, 4, 5, 6, 7, 8], offering potential solutions for sustainable clean energy. To date, however, large-scale TE applications remain limited due to the relatively low energy conversion efficiency of known materials. Improving the efficiency and finding suitable TE materials that function at different temperatures remain important tasks in materials science research.
The TE performance can be quantified by the figure of merit value $ZT=\frac{S^{2}\sigma}{\kappa T}$, where $S$ is the Seebeck coefficient (the induced voltage in response to a temperature gradient), $\sigma$ is the electrical conductivity, $\kappa$ is the thermal conductivity, and $T$ is the temperature. One approach to enhance $ZT$ is by increasing the power factor ($S^{2}\sigma$), via band engineering of carrier concentration and mobility, among other factors [9, 10, 11, 12, 13, 14, 15]. Another approach is to find materials with low thermal conductivity [16, 17, 18].
[…]
Computational materials modeling has played an important role in providing predictions and critical insights into the thermal conducting behavior of materials [27, 28, 29, 30, 31, 32]. Traditionally, density functional theory (DFT) is the standard computational workforce for accurate calculations of $\kappa_{L}$ from first principles. However, its relatively high computational cost limits large-scale investigations of $\kappa_{L}$ in new materials. More recently, data-driven machine learning (ML) approaches have become popular and powerful tools for materials modeling and discovery [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. This popularity and improvement in ML research largely result from advancements in computer architectures and ML algorithms, as well as from the increasing availability of open materials databases. ML algorithms can learn from training data by identifying connections through linear or non-linear relationships between target properties and input features. Once trained, ML models can achieve highly efficient and often accurate large-scale predictions.
In this study, we utilize combined machine learning (ML) predictions and density functional theory (DFT) calculations to discover materials with low lattice thermal conductivity. Specifically, we develop ensemble-tree ML models with input features based on the chemical formula and atomic configurations to quickly estimate $\kappa_{L}$ of a given material. For promising low-$\kappa$ materials identified by our ML models, we further validate the results by performing DFT calculations to evaluate $\kappa_{L}$ directly from first principles. In particular, we find that the chemical compositions A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) of orthorhombic crystal symmetry exhibit ultra-low lattice thermal conductivity ($\kappa_{L}\sim 0.1-1.0$ W/mK). Our DFT calculations of the transport and thermoelectric properties further indicate that some of these materials like K2CdPb can exhibit a $ZT\geq 1.0$ near room temperature, which are thereby promising for low-temperature thermoelectric application [46].
The rest of the paper is organized as follows: Section II presents the computational details of machine learning (ML) models and first-principles density functional theory (DFT) calculations. Section III presents the ML and DFT predictions of low-$\kappa_{L}$ materials and their thermoelectric properties. Finally, Section IV concludes the paper by summarizing our main findings.
II COMPUTATIONAL METHODS
II.1 Machine Learning Simulation
Data Acquisition and Feature Creation – Our machine learning (ML) models aim at predicting the target property of lattice thermal conductivity $\kappa_{L}$ for a given compound. The ML training dataset is sourced from the TE Design Lab, which is a virtual platform hosting a database of calculated thermoelectric properties [47]. From this database, we select a total of 1900 compounds with $\kappa_{L}$ in the range of $0-1100$ W/mK. For all compounds in the selected dataset, we then use the Matminer package [48] to generate 61 input features based on their chemical formula and atomic configurations [49]. These features can be broadly categorized as structural features and elemental features. Specifically, 7 structural features include the space group, volume per atom, packing fraction, unit-cell density, bond length, bond angle, and cohesive energy. Moreover, 18 elemental features include the atomic mass, atomic radius, atomic number, periodic table group, row number, block number, Mendeleev number, molar volume, boiling point, melting temperature, Pauling electronegativity, first ionization energy, covalent radius, and volume per atom from ground state, as well as the average number of $s$, $p$, $d$, and $f$ valence electrons. Since our dataset contains compounds ranging from unary to quinary materials, each elemental feature can be expanded by calculating the minimum, maximum, and weighted average of the constituent chemicals, resulting in a total of 54 $(=18\times 3)$ elemental features. Overall, 61 $(=7+54)$ features are used in training the ML models.
Model Training and Validation – Our supervised ML tasks utilize Random Forests as the underlying algorithm [50, 51]. Random Forest is an ensemble method consisting of multiple decision trees. Each tree is trained on a randomly selected subset of features and samples. The Random Forest algorithm then averages the results of all trees to make the final prediction, which generally reduces the overfitting problem associated with a single decision tree. Random Forest ML models are relatively easy to train and often produce highly accurate results. To further reduce overfitting, we also pre-prune the trees by limiting their depth. Specifically, we use 90% of our input data as the training-validation set and apply the GridSearchCV technique from the scikit-learn library [52] to determine the optimal tree depth via 10-fold cross-validation. The remaining 10% of the input data serves as the unbiased test set to evaluate the final model performance. After training and evaluation, we then use the ML model to predict the lattice thermal conductivity $\kappa_{L}$.
II.2 First-Principles Calculation
For promising low-$\kappa_{L}$ materials identified by ML models, we further perform first-principles density functional theory (DFT) calculations to validate their thermoelectric properties. Our calculations are based on the Vienna Ab initio Simulation Package (VASP, version 5.4.4) [53, 54], which is a highly efficient and accurate plane-wave pseudopotential DFT code. We adopt the projector-augmented-wave (PAW) potentials [55, 56] and utilize the Perdew-Burke-Ernzerhof generalized gradient approximation (GGA-PBE) functional [57]. The plane-wave cutoff energy is set to 500 eV, and a fine $\Gamma$-centered Monkhorst-Pack grid of $19\times 19\times 19$ points is used for Brillouin zone integration [58]. For a given crystal structure, we first fully relax the lattice parameters and atomic positions. The convergence criteria for the electronic and ionic relaxation loops are set to $10^{-8}$ eV per unit cell and $10^{-4}$ eV/Å, respectively.
After structure relaxation, we compute the thermoelectric properties ($S$, $\sigma$, and $\kappa_{e}$) using the BoltzTraP2 package [59], which is based on Boltzmann transport theory with a constant relaxation time approximation. The lattice thermal conductivity ($\kappa_{L}$) is obtained through first-principles phonon calculations using the Phonopy and Phono3py [60, 61] packages, which are based on finite-displacement supercell approaches. Phonopy computes the phonon spectra at the harmonic or quasi-harmonic level. Phono3py evaluates phonon-phonon interactions and $\kappa_{L}$ from the Peierls-Boltzmann equation [62]. In the supercell calculations, the atomic displacement is set to 0.02 Å, and the real-space interaction cutoff distance is set to 4.0 Å. For the second-order (harmonic) and third-order (anharmonic) phonon calculations, $3\times 3\times 3$ supercells with a $5\times 5\times 5$ k-mesh and $2\times 2\times 2$ supercells with a $9\times 9\times 9$ k-mesh are employed, respectively. A phonon q-point sampling mesh of $21\times 21\times 21$ points is used. The theoretical crystal structure in this study is visualized using the VESTA software (version 3.4.8) [63].
Figure 1: (a) Histogram for the 1900 training samples of lattice thermal conductivity ($\kappa_{L}$) selected from the TE Design Lab [47]. The distribution spans nearly five orders of magnitude and is plotted on a logarithmic scale. (b) False-color intensity plot showing the frequency of each element in the 1900 training dataset. Elements not present in the list are shown in gray. The figure was created using the open-source software Periodic Trend Plotter [64].
III RESULTS AND DISCUSSION
Figure 1(a) shows the distribution of $\kappa_{L}$ for the 1900 compounds in our training dataset from the TE Design Lab [47]. Since the range of the distribution spans nearly 5 orders of magnitude, it is plotted on a logarithmic scale. Eventually, ML models are trained to predict $\log(\kappa_{L})$. For the accuracy and generalizability of our ML models, we ensure that our dataset is diverse in chemical composition (from unary to quinary compounds) and crystal structure (with 140 different space groups). In particular, among the 1900 samples, 7 are unary, 418 are binary, 1143 are ternary, 328 are quaternary, and 4 are quinary. These compounds contain 61 different atomic elements; the frequency each element appears in the compound list is represented by the false-color intensity plot in Figure 1(b). Gray color means that the element is not present.
As discussed in Section II, our ML models are based on Random Forests trained with 61 features [49] generated by the Matminer package [48]. The coefficient of determination $R^{2}$ is used to evaluate the model performance:
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y_{i}})^{2}}{\sum_{i}(y_{i}-\overline{y})^{2}},$ (1)
where $y_{i}$, $\hat{y_{i}}$, $\overline{y}$ are the actual value (for the $i$-th entry), the predicted value, and the mean of the actual values, respectively. $R^{2}$ ranges from 0 to 1, with $R^{2}=1$ indicating a perfect prediction. Figure 2 shows the resulting ML model performance on predicting $\log(\kappa_{L})$. The blue and yellow circles represent data from the training-validation set (90%) and the test set (10%), respectively. A red dashed line is also plotted as a guide to the ideal line where the predicted values match the actual values. Our model achieves an $R^{2}=0.96$ for the training-validation set and $R^{2}=0.88$ for the test set, indicating that our ML model provides a fairly accurate prediction of $\log(\kappa_{L})$.
Figure 2: Evaluation of the Random Forests model in predicting the logarithmic value of lattice thermal conductivity, $\log(\kappa_{L})$. The training and test sets consist of 90% and 10% of our total dataset (1900 samples), respectively. When the machine learning prediction perfectly matches the actual value, the data point will fall on the red dashed line. The model achieves relatively high $R^{2}$ scores of 0.96 and 0.88 for the training and test sets, respectively.
Random Forest models also provide information on feature importance in the ML predictions. Among the features under study, the atomic bond length is found to be the most significant factor affecting $\kappa_{L}$. In an ideal gas model, lattice thermal conductivity is approximated as
$\kappa_{L}=\frac{1}{3}v^{2}_{s}c_{v}\tau_{s},$ (2)
where $v_{s}$ is the phonon velocity, $c_{v}$ is the specific heat, and $\tau_{s}$ is the phonon relaxation time. Among these three parameters, the phonon relaxation time is related to bond-strength anharmonicity [65, 66, 67], which is correlated with bond length. In particular, a longer bond length is prone to causing anharmonic vibrations, as the interatomic force constant decreases with increasing bond length. Anharmonicity then facilitates collisions between different phonon modes. As anharmonicity increases, the phonon relaxation time decreases, which in turn leads to a reduction in lattice thermal conductivity.
We note that the strength of anharmonicity also can be evaluated by the Grüneisen parameter:
$\gamma=\frac{V}{\omega}\frac{\partial\omega}{\partial V},$ (3)
where $V$ is the crystal volume and $\omega$ is the phonon frequency. Within the harmonic approximation, the thermal expansion is zero on average. In the presence of anharmonicity, the phonon frequency can vary as the volume changes with temperature. Therefore, a larger Grüneisen parameter indicates stronger anharmonicity and a lower lattice thermal conductivity. In fact, based on the Debye-Callaway model [68, 69], the lattice thermal conductivity can be approximately evaluated as
$\kappa_{L}\approx\frac{Mv^{3}_{m}}{TV^{2/3}\gamma^{2}}\frac{1}{N^{1/3}},$ (4)
where $M$, $v_{m}$, $T$, $V$, $\gamma$, and $N$ represent the average mass, the mean speed of sound, the temperature, the average volume per atom, the Grüneisen parameter, and the number of atoms per primitive unit cell, respectively. The above formula shows that $\kappa_{L}$ is inversely proportional to $\gamma^{2}$ and $V^{2/3}$. Indeed, in addition to bond length, the volume per atom is evaluated by our ML models as the second most important feature affecting $\kappa_{L}$. Overall, the feature importances align well with the above approximated models for $\kappa_{L}$, demonstrating that our ML models are reasonable and adequate.
We next apply the ML models to predict materials with low $\kappa_{L}$. Recently, Zintl-phase compounds have attracted significant attention due to their strong anharmonic properties, which could lead to low lattice thermal conductivities [70, 71, 72, 73, 74, 75, 76, 77]. The Zintl phase refers to compounds formed by alkali metals (group I) or alkaline-earth metals (group II) combined with $p$-block metals or metalloids (from groups III-VI). Other recent studies have also shown that Zintl-phase compounds can achieve ultra-low $\kappa_{L}$ by introducing a heavy element, cadmium (Cd) [78, 20]. For these reasons, we focus on applying our ML models to Cd-based Zintl-phase materials. Specifically, we consider A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) with orthorhombic symmetry and space group Ama2 (No. 40) [79]; Figure 3(a) shows the corresponding crystal structure for K2CdPb. As seen in Table I, the $\kappa_{L}$ values predicted by our ML models for the nine compositions A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) range from $0.69$ to $0.95$ W/mK, indicating that all these compounds are potential low-$\kappa_{L}$ materials.
Figure 3: (a) Primitive-cell crystal structure of K2CdPb, with orthorhombic symmetry and space group Ama2 (No. 40). (b) Lattice thermal conductivity $\kappa_{L}$ of K2CdPb as a function of the neighbor interaction cutoff distance in the temperature range of $300-1200$ K. (c) $\kappa_{L}$ of K2CdPb as a function of temperature for different cutoff distances. (d) $\kappa_{L}$ computed with a cutoff distance of 4 Å for nine different Cd-compounds.
To validate the ML predictions, we further perform first-principles calculations to directly compute $\kappa_{L}$ and other thermoelectric properties for the nine compounds under study. We first focus on K2CdPb [79], whose primitive cell structure is shown in Figure 3(a). To ensure accurate calculation of $\kappa_{L}$, we conduct a convergence test for the neighbor interaction cutoff distance. Figures 3(b) and 3(c) show the convergence tests for K2CdPb as functions of temperature and cutoff distance. Notably, the $\kappa_{L}$ computed with a cutoff of 4 Å is very close to the result obtained with a 5 Å cutoff. Therefore, for both accuracy and efficiency considerations, we have adopted a cutoff distance of 4 Å for the other compounds. Figure 3(d) shows the computed $\kappa_{L}$ as a function of temperature for the nine compounds A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge). Near room temperature, all compounds exhibit a $\kappa_{L}$ below 1.0 W/mK, in very good agreement with our ML predictions. As the temperature increases, $\kappa_{L}$ further decreases as more phonons are excited and cause additional phonon scattering, leading to a reduction in $\kappa_{L}$. These results reveal that the nine compounds under study are all low-$\kappa_{L}$ materials for potential thermal management and insulation applications.
Methods K2CdPb K2CdSn K2CdGe Na2CdPb Na2CdSn Na2CdGe Li2CdPb Li2 CdSn Li2CdGe
ML 0.69 0.79 0.8 0.84 0.76 0.87 0.95 0.71 0.77
DFT 0.295 0.346 0.323 0.022 0.119 0.086 0.032 0.215 0.028
ML + Weight 5 0.56 0.35 0.37 0.45 0.63 0.119 0.68 0.27 0.4
ML + Weight 10 0.469 0.29 0.62 0.43 0.37 0.58 0.76 0.4 0.22
Table 1: Machine learning (ML) and density functional theory (DFT) predictions of the lattice thermal conductivity $\kappa_{L}$ (in units of W/mK) for different Cd-compounds. The ML models are based on Random Forests. The terms “ML + Weight 5” and “ML + Weight 10” indicate a weighting factor of 5 and 10, respectively, on samples with $\log(\kappa_{L})\leq 0$ when training the ML models, which places more weight on the low-$\kappa_{L}$ materials.
Before discussing other thermoelectric properties, we address the small discrepancies between the ML and DFT results in Table I. First, we note that the ML models were trained to predict $\log(\kappa_{L})$ rather than $\kappa_{L}$ itself. Therefore, a small error in the logarithmic value can be amplified in the actual value. Second, the ML prediction in Table I is generally slightly larger than the DFT calculation. One reason for this discrepancy is likely due to the training data distribution. Specifically, while we are interested in discovering materials with low thermal conductivity (i.e., $\log(\kappa_{L})\leq 0$), most of the training data in Figure 1(a) exhibit $\log(\kappa_{L})\geq 0$. Therefore, one could potentially enhance the ML prediction accuracy by training a weighted ML model. Indeed, when we train additional ML models by weighting the training samples with $\log(\kappa_{L})\leq 0$ by a factor of 5 to 10, the predicted $\kappa_{L}$ values become smaller and align more closely with the DFT results, as seen in Table I. Notably, the performance of the ML model with a weight factor of 10 is not better than that with a weight factor of 5. Therefore, the sample weight factor has an optimal range and cannot be increased indefinitely. Finally, we note that some ML predictions from the weighted ML models remain larger than the DFT results (e.g., for the sodium compounds in Table I). This discrepancy may be attributed to the fact that tree models only interpolate so cannot predict values beyond the range of the training dataset. The apparent, albeit small, differences between the ML and DFT predictions are likely associated with the above factors.
Figure 4: Thermoelectric properties of K2CdPb from first-principles calculations: (a) Seebeck coefficient ($S$), (b) electrical conductivity divided by the relaxation time ($\sigma$/$\tau$), (c) electronic thermal conductivity divided by the relaxation time ($\kappa_{e}$/$\tau$), as a function of the carrier concentration $n$ (in log scale) over the temperature range of $300-1200$ K.
We next turn our attention to the thermoelectric properties. Figure 4 displays the DFT calculations for K2CdPb: Seebeck coefficient $S$ [panel (a)], electrical conductivity divided by the relaxation time $\sigma/\tau$ [panel (b)], and electronic thermal conductivity divided by the relaxation time $\kappa_{e}/\tau$ [panel (c)], as a function of the carrier concentration $n$ in the temperature range $300-1200$ K. In general, the Seebeck coefficient exhibits a more complex temperature and carrier concentration dependence, but its behavior can be understood qualitatively by considering that of a simple parabolic band [80, 81]:
$S=\frac{8\pi^{2}k^{2}_{B}}{3eh^{2}}m^{*}T(\frac{\pi}{3n})^{2/3}.$ (5)
Here, $k_{B}$, $e$, $h$, and $m^{*}$ are the Boltzmann constant, electron charge, Plank constant, and carrier effective mass, respectively. Equation (5) dictates that a higher temperature $T$ or a lower carrier concentration $n$ would result in a larger Seebeck coefficient $S$. These $T$ and $n$ dependences are indeed consistent with those shown in Figure 4(a), especially in the high carrier concentration regime ($n>10^{20}$ cm-3). In contrast, the low concentration regime exhibits an opposite trend, where $S$ is reduced at higher temperatures. This anomalous behavior is caused by the bipolar effect [82, 83, 84], where thermal excitations generate both electrons and holes, which contribute opposite signs and lead to an overall reduced $S$.
[…]
Finally, since low-$\kappa$ materials can be good candidates for thermoelectrics, we also compute their figure of merit, $ZT=\frac{S^{2}\sigma}{\kappa T}$, where the thermal conductivity $\kappa=\kappa_{e}+\kappa_{L}$ includes both electronic and lattice contributions. Figures 5(a)-(c) show the $ZT$ values respectively for K2CdPb, K2CdSn, and K2CdGe, as functions of carrier concentration $n$ (in log scale) and temperature $T$. In all three compounds, the $ZT$ values can exceed 1.0. In particular, K2CdPb exhibits a peak $ZT$ value of $\sim 1.4$ at $n=2\times 10^{20}$ cm-3 and $T=400$ K, making it a promising low-temperature thermoelectric material. In contrast, K2CdSn and K2CdGe show peak $ZT$ values of $\sim 1.1$ near $n=9\times 10^{20}$ cm-3 and $T=900$ K, and they are more suitable for thermoelectric applications at higher temperatures. For the other compounds based on sodium and lithium listed in Table I, the $ZT$ values are less than 1.0, making them unsuitable for practical thermoelectric applications. However, they could still be potential candidates for thermal insulation materials.
Figure 5: $ZT$ values of thermoelectric performance from first-principles calculations for (a) K2CdPb, (b) K2CdSn, and (c) K2CdGe, as a function of the carrier concentration $n$ (in log scale) over the temperature range of $300-1200$ K.
IV CONCLUSIONS
We have developed machine learning (ML) models using Random Forests to efficiently predict the lattice thermal conductivity ($\kappa_{L}$) of a given chemical compound. We have also conducted first-principles density functional theory (DFT) calculations to validate the ML predictions. The results indicate that the nine Zintl-phase Cd-compounds A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) with orthorhombic crystal symmetry all exhibit very low lattice thermal conductivities, with $\kappa_{L}\leq 1.0$ W/mK. Our DFT calculations of the figure of merit, $ZT$, for thermoelectric performance further showed that K2CdPb exhibits a peak $ZT\sim 1.4$ near 400 Kelvin, making it a promising low-temperature thermoelectric material. Additionally, K2CdSn and K2CdGe were found to display $ZT$ values of $\sim 1.1$ at 900 Kelvin, suggesting they could be candidate thermoelectrics at higher temperatures.
For Li2CdX and Na2CdX (X = Pb, Sn, and Ge), the $ZT$ values are less than 1.0, indicating more limited practical thermoelectric applications. Nevertheless, their ultra-low lattice thermal conductivities make these materials potentially useful for thermal management and insulation applications. Overall, our study demonstrated that data-driven ML methods are powerful tools for large-scale materials modeling and discovery. Experimental verification of our ML and DFT predictions on the thermoelectric properties of the Zintl-phase Cd-compounds would be an important next step. Further theoretical exploration of additional low-$\kappa_{L}$ and high-$ZT$ materials using a combined ML and DFT methodology will continue to be an important area of future research.
Acknowledgements
This work is supported by the U.S. Air Force Office of Scientific Research (AFOSR) under Award No. FA2386-21-1-4060. The calculations were performed on the Frontera supercomputer at the Texas Advanced Computing Center. Frontera is made possible by National Science Foundation Award No. OAC-1818253.
Data Availability Statement
The data for training machine learning (ML) models and the resulting ML code for predicting lattice thermal conductivity can be found online at the following weblink: https://github.com/CMLUAB/ML_lattice-themal-conductivity. The data for first-principles calculations are available upon request from the authors.

### v2

@@
 Improving the efficiency and finding suitable TE materials that function at different temperatures remain important tasks in materials science research.
-The TE performance can be quantified by the figure of merit value $ZT=\frac{S^{2}\sigma}{\kappa T}$, where $S$ is the Seebeck coefficient (the induced voltage in response to a temperature gradient), $\sigma$ is the electrical conductivity, $\kappa$ is the thermal conductivity, and $T$ is the temperature.
+The TE performance can be quantified by the figure of merit value $ZT=S^{2}\sigma T/\kappa$, where $S$ is the Seebeck coefficient (the induced voltage in response to a temperature gradient), $\sigma$ is the electrical conductivity, $T$ is the temperature, and $\kappa$ is the thermal conductivity.
 One approach to enhance $ZT$ is by increasing the power factor ($S^{2}\sigma$), via band engineering of carrier concentration and mobility, among other factors [9, 10, 11, 12, 13, 14, 15].
@@
 Overall, 61 $(=7+54)$ features are used in training the ML models.
+We note that several features, such as average atomic mass and volume (which is related to atomic radius), are relevant parameters for estimating $\kappa_{L}$ in known empirical formulas [50, 51].
+There, $\kappa_{L}$ is also expected to be proportional to the mean sound velocity $v_{m}$ (or the Debye temperature $\Theta_{D}$) cubed.
+It was shown that ML models can accurately predict $v_{m}$ and $\Theta_{D}$ using features simply derived from chemical compositions and crystal symmetry [52].
+Therefore, it is anticipated that ML models trained here with the 61 features under study can perform well in predicting $\kappa_{L}$ [49].
 Model Training and Validation – Our supervised ML tasks utilize Random Forests as the underlying algorithm [53, 54].
@@
 The theoretical crystal structure in this study is visualized using the VESTA software (version 3.4.8) [66].
+III Results and Discussion
 Figure 1: (a) Histogram for the 1900 training samples of lattice thermal conductivity ($\kappa_{L}$) selected from the TE Design Lab [47].
-The distribution spans nearly five orders of magnitude and is plotted on a logarithmic scale.
+The distribution spans nearly five orders of magnitude and is plotted on a base-10 logarithmic scale.
 (b) False-color intensity plot showing the frequency of each element in the 1900 training dataset.
@@
 The figure was created using the open-source software Periodic Trend Plotter [67].
-III RESULTS AND DISCUSSION
 Figure 1(a) shows the distribution of $\kappa_{L}$ for the 1900 compounds in our training dataset from the TE Design Lab [47].
-Since the range of the distribution spans nearly 5 orders of magnitude, it is plotted on a logarithmic scale.
+Since the range of the distribution spans nearly 5 orders of magnitude, it is plotted on a base-10 logarithmic scale.
 Eventually, ML models are trained to predict $\log(\kappa_{L})$.
@@
 40) [80]; Figure 3(a) shows the corresponding crystal structure for K2CdPb.
-As seen in Table I, the $\kappa_{L}$ values predicted by our ML models for the nine compositions A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) range from $0.69$ to $0.95$ W/mK, indicating that all these compounds are potential low-$\kappa_{L}$ materials.
+As seen in Table 1, the $\kappa_{L}$ values predicted by our ML models for the nine compositions A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) range from $0.69$ to $0.95$ W/mK, indicating that all these compounds are potential low-$\kappa_{L}$ materials.
 Figure 3: (a) Primitive-cell crystal structure of K2CdPb, with orthorhombic symmetry and space group Ama2 (No.
 40).
-(b) Lattice thermal conductivity $\kappa_{L}$ of K2CdPb as a function of the neighbor interaction cutoff distance in the temperature range of $300-1200$ K.
-(c) $\kappa_{L}$ of K2CdPb as a function of temperature for different cutoff distances.
+(a) Lattice thermal conductivity $\kappa_{L}$ of K2CdPb as a function of the neighbor interaction cutoff distance in the temperature range of $300-1200$ K.
+(a) $\kappa_{L}$ of K2CdPb as a function of temperature for different cutoff distances.
 (d) $\kappa_{L}$ computed with a cutoff distance of 4 Å for nine different Cd-compounds.
+Methods K2CdPb K2CdSn K2CdGe Na2CdPb Na2CdSn Na2CdGe Li2CdPb Li2 CdSn Li2CdGe
+ML 0.69 0.79 0.8 0.84 0.76 0.87 0.95 0.71 0.77
+DFT 0.295 0.346 0.323 0.022 0.119 0.086 0.032 0.215 0.028
+ML + Weight 5 0.56 0.35 0.37 0.45 0.63 0.119 0.68 0.27 0.4
+ML + Weight 10 0.469 0.29 0.62 0.43 0.37 0.58 0.76 0.4 0.22
+Table 1: Machine learning (ML) and density functional theory (DFT) predictions of the lattice thermal conductivity $\kappa_{L}$ (in units of W/mK) for different Cd-compounds.
+The ML models are based on Random Forests.
+The terms “ML + Weight 5” and “ML + Weight 10” indicate a weighting factor of 5 and 10, respectively, on samples with $\log(\kappa_{L})\leq 0$ when training the ML models, which places more weight on the low-$\kappa_{L}$ materials.
 To validate the ML predictions, we further perform first-principles calculations to directly compute $\kappa_{L}$ and other thermoelectric properties for the nine compounds under study.
@@
 These results reveal that the nine compounds under study are all low-$\kappa_{L}$ materials for potential thermal management and insulation applications.
-Methods K2CdPb K2CdSn K2CdGe Na2CdPb Na2CdSn Na2CdGe Li2CdPb Li2 CdSn Li2CdGe
-ML 0.69 0.79 0.8 0.84 0.76 0.87 0.95 0.71 0.77
-DFT 0.295 0.346 0.323 0.022 0.119 0.086 0.032 0.215 0.028
-ML + Weight 5 0.56 0.35 0.37 0.45 0.63 0.119 0.68 0.27 0.4
-ML + Weight 10 0.469 0.29 0.62 0.43 0.37 0.58 0.76 0.4 0.22
-Table 1: Machine learning (ML) and density functional theory (DFT) predictions of the lattice thermal conductivity $\kappa_{L}$ (in units of W/mK) for different Cd-compounds.
-The ML models are based on Random Forests.
-The terms “ML + Weight 5” and “ML + Weight 10” indicate a weighting factor of 5 and 10, respectively, on samples with $\log(\kappa_{L})\leq 0$ when training the ML models, which places more weight on the low-$\kappa_{L}$ materials.
-Before discussing other thermoelectric properties, we address the small discrepancies between the ML and DFT results in Table I.
+We further note that our reported $\kappa_{L}\sim 0.3$ W/mK for K2CdPb and K2CdSn is consistent with previous DFT studies [80].
+Meanwhile, we also find that K2CdGe exhibits a comparable theoretical $\kappa_{L}\sim 0.3$ W/mK.
+This result is not surprising, given the chemical similarity of the Pb, Sn, and Ge elements.
+Likewise, for the Li- and Na-based compounds explored here, while our DFT calculations may underestimate their $\kappa_{L}$ values, they are also anticipated to be low lattice thermal conductivity materials, based on their chemical similarity to the K-based compounds.
+In fact, depending on carrier concentration and the underlying temperature, materials with ultra-low $\kappa_{L}\sim 0.1$ W/mK or lower have been reported in the literature [81, 82, 83, 84].
+It would be an important future task to verify our predictions both theoretically (e.g., with different DFT functionals and supercell sizes) and experimentally, through potential synthesis and characterization of the proposed materials.
+Before discussing other thermoelectric properties, we address the small discrepancies between the ML and DFT results in Table 1.
 First, we note that the ML models were trained to predict $\log(\kappa_{L})$ rather than $\kappa_{L}$ itself.
@@
 […]
-Finally, since low-$\kappa$ materials can be good candidates for thermoelectrics, we also compute their figure of merit, $ZT=\frac{S^{2}\sigma}{\kappa T}$, where the thermal conductivity $\kappa=\kappa_{e}+\kappa_{L}$ includes both electronic and lattice contributions.
+Finally, since low-$\kappa$ materials can be good candidates for thermoelectrics, we also compute their figure of merit, $ZT=S^{2}\sigma T/\kappa$, where the thermal conductivity $\kappa=\kappa_{e}+\kappa_{L}$ includes both electronic and lattice contributions.
 Figures 5(a)-(c) show the $ZT$ values respectively for K2CdPb, K2CdSn, and K2CdGe, as functions of carrier concentration $n$ (in log scale) and temperature $T$.
 In all three compounds, the $ZT$ values can exceed 1.0.
-In particular, K2CdPb exhibits a peak $ZT$ value of $\sim 1.4$ at $n=2\times 10^{20}$ cm-3 and $T=400$ K, making it a promising low-temperature thermoelectric material.
+As an example to estimate the $ZT$ value, for K2CdPb at $T=400$ K and $n=2\times 10^{20}$ cm-3, the relevant parameters from our calculations are $S\sim 1.8\times 10^{-4}$ V/K, $\sigma\sim 0.1\times 10^{5}$ 1/$\Omega$m, and $\kappa=\kappa_{e}+\kappa_{L}\sim 1.0$ W/mK.
+Together, these values lead to a figure of merit $ZT=S^{2}\sigma T/\kappa(\sim 3.24\times 10^{-8}\times 0.1\times 10^{5}\times 400/1.0)\sim 1.3-1.4$ for K2CdPb, making it a promising low-temperature thermoelectric material.
 In contrast, K2CdSn and K2CdGe show peak $ZT$ values of $\sim 1.1$ near $n=9\times 10^{20}$ cm-3 and $T=900$ K, and they are more suitable for thermoelectric applications at higher temperatures.
@@
 Figure 5: $ZT$ values of thermoelectric performance from first-principles calculations for (a) K2CdPb, (b) K2CdSn, and (c) K2CdGe, as a function of the carrier concentration $n$ (in log scale) over the temperature range of $300-1200$ K.
-IV CONCLUSIONS
+IV Conclusion
 We have developed machine learning (ML) models using Random Forests to efficiently predict the lattice thermal conductivity ($\kappa_{L}$) of a given chemical compound.